OpenStudy (anonymous):

y=x^2+2x-3 x-intercepts

OpenStudy (anonymous):

x-intercepts occur where y=0. Set y=0 and solve for x: \[0=x^2+2x-3\] The right side will factor: \[0=(x+3)(x-1)\] Okay from there?

OpenStudy (anonymous):

x-intercepts are where y=0 \[0=(x+3)(x-1)\] so intercepts are at: \[x=-3, x=1\]